please show me how to do this problem use the formula y =20/x to find y when x =10
Answer by TimothyLamb(4379)

y = 20/x
---
when x = 10 ...
y = 20/x
y = 20/10
y = 2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php